# Quantifying task-relevant representational similarity using decision variable correlation

### v1

@@ C.1 General information on the DVC framework
The DVC method is computationally efficient and stable, as dimensionality reduction is applied before attempting to decode the DV using LDA. All experiments were performed on Intel(R) Core i7-14700K CPU without resorting to GPU usage. Computing DVC between a pair of models take  30 seconds on average, with the total compute rounding to  25 hours.
@@ C.2 Verification of the robustness of DVC results
While the guiding principle behind DVC is general and intuitive, specific implementation choices carry implications on the numerical stability and robustness to different data distributions. We thus experimented with different algorithms and hyperparameter choices and found that they do not affect the main conclusions drawn in this study. Specifically, we want to verify if the choice of PC dimensions might change the conclusions in this study. First we note that with 25 PC dimensions (for each split), all the monkey recordings and model representations achieve high binary linear seprability (Figure A.2a), and that 8-way logistic regression accuracy plateaus early on (Figure A.2b). In addition, the main result is robust with varying PC dimensions (Figure A.3a,3b).
@@
Other choices also do not affect the results. Pearson’s correlation is easily biased by extreme values. Although the DVs appear normally distributed, we substituted Pearson’s correlation with Spearman rank correlation and found that the trend persists (Figure A.3c). In addition, we tested shrink regularization for LDA, which could enhance stability of the procedure, and find that the result is robust to the choice of LDA solver (Figure A.3d).
@@
Figure A.2: Dimensionality reduction retains task-relevant information. a. LDA score of all monkeys (red) and networks are high. b. As dimensionality increases, decoding performance using logistic regression plateaus.
Figure A.3: The main result is robust to the following choices: a. PCA dim =10; b. PCA dim = 50; c. Correlation measures (Pearson vs. Spearman); d. LDA solvers (SVD vs. eigen+shrinkage)
C.3 Split normalization under uncommon conditions
@@
While essential to this method, the split normalization procedure could behave in unexpected ways. For example, it could result in a DVC value larger than 1 when the internal noise is high. None of the normalized correlations between different representations reached the cap. In addition, we took the absolute value before calculating the geometric means. When there is no correlation between two representations, this would bias the normalized DVC to a small positive value. None of the representations in this study fall in this range.
@@
Appendix D More on the effect of behavioral decoding on Cohen’s Kappa
@@
Geirhos et al. have discussed an important caveat of applying Cohen’s Kappa in details: Cohen’s Kappa is bounded by the accuracies of the subjects. Intuitively, the accuracy difference $|p_{i}-p_{j}|$ provides an upper bound on kappa. When the accuracy difference is high, one subject is often right, while the other is often wrong, then their behaviors cannot be consistent. From the original behavioral data, it is clear that human subjects perform with high accuracy but models perform poorly (Figure A.4a). While the authors used simulations to show that under the condition that the subjects act independently, accuracy difference is not necessarily correlated with model performance, it seems that the low model-human consistency is a direct result of their accuracy difference (Figure A.4b,4c).
Figure A.4: Recapitulation of results from Geirhos ’edges’ behavioral data. a. Accuracies of human subjects (red) are generally a lot higher than the models tested. b. Cohen’s Kappa between models and human subjects are low while Cohen’s Kappa between models and between human subjects are high, consistent with the original report. c. Cohen’s Kappa bounded by accuracy difference $|p_{i}-p_{j}|$.
@@
Further, we want to verify if the high model-model consistency reported in the original work might be inflated by the low-accuracy high-bias condition caused by the choice of decoder (directly aggregating the probabilities). To test this, we take the original stimuli provided by Geirhos et al., and calculate the Cohen’s Kappa between the models by a) taking the average of the probability of the corresponding ImageNet-1k subclasses or b) training a 5-fold cross-validated logistic regression decoder on the representations in the penultimate layer. The result shows that compared to a), b) achieves higher accuracy (Figure A.5b, 5c), exhibits less bias towards certain categories (Figure A.5d,5e) and results in significantly lower model-model consistency as measured by Cohen’s Kappa (Figure A.5a).
@@
Figure A.5: Using logistic regression decoder results in higher accuracy, lower bias and lower Cohen’s Kappa estimate compared to mean probability decoder, estimated on the ’edges’ images. a. Cohen’s Kappa estimated using a mean probability decoder is significantly higher than that estimated by a logistic regression decoder. b. Behavioral accuracy of mean probability decoder. c. Behavioral accuracy of logistic regression decoder. d. Choice histogram of the mean probabilities decoder. e. Choice histogram of the logistic regression decoder.
@@
While the shared behavioral bias that results from aggregating probabilities from the original ImageNet-1k classes is very interesting, it does make the human-model consistency and the model-model consistency a lot more ambiguous. Therefore we think it is more suitable to use a stronger behavioral decoder or to use DVC when applicable.
Appendix E Broader Impacts
@@
We expect DVC to be a broadly applicable approach to study the similarity of the brain and neural network models. On the positive side, the method and the results discussed in this paper could redirect community focus away from brute-force scaling and toward more targeted investigations into task-relevant representation alignment and model-brain convergence. It could in the long term lead to models that are more brain-like, thus greatly facilitating research in fields like neuroscience, cognitive science and AI interpretability and safety. However, while DVC offers a biologically grounded lens for comparing model and brain representations, promoting alignment with biological brains might inadvertently constrain models in certain domains where brain cognition is suboptimal.

### v3

@@ C.1 General information on the DVC framework
The DVC method is computationally efficient and stable, as dimensionality reduction is applied before attempting to decode the DV using LDA. All experiments were performed on Intel(R) Core i7-14700K CPU without resorting to GPU usage. Computing DVC between a pair of models takes  30 seconds on average, with the total compute rounding to  30 hours.
@@ C.2 Verification of the robustness of DVC results
While the guiding principle behind DVC is general and intuitive, specific implementation choices carry implications on the numerical stability and robustness to different data distributions. We thus experimented with different algorithms and hyperparameter choices and found that they do not affect the main conclusions drawn in this study. Specifically, we want to verify if the choice of PC dimensions might change the conclusions in this study. First we note that with 25 PC dimensions (for each split), all the monkey recordings and model representations achieve high binary linear seprability (Fig. A.2a), and that 8-way logistic regression accuracy plateaus early on (Fig. A.2b). In addition, the main result is robust with varying PC dimensions (Fig. A.3a,b).
@@
Other choices also do not affect the results. Pearson’s correlation is easily biased by extreme values. Although the DVs appear normally distributed, we substituted Pearson’s correlation with Spearman rank correlation and found that the trend persists (Fig. A.3c). In addition, we tested shrink regularization for LDA, which could enhance stability of the procedure, and find that the result is robust to the choice of LDA solver (Fig. A.3d).
@@
Figure A.2: Dimensionality reduction retains task-relevant information. (a) LDA score of all monkeys (red) and networks are high. (b) As the dimensionality increases, decoding performance using logistic regression plateaus.
C.3 Consistency of the DVC results across different linear decoders
The choice of using LDA as a decoder is not coincidental. Under the assumption of Gaussian distribution and in the setting of pairwise classification, a linear decoder provides the best classifier. We calculated DVC between the two monkey subjects using logistic regression and linear support vector machine (SVM). We find that the DVC is highly consistence across different linear decoders. In addition, across repeats (random split normalization), DVC estimation varied little. On the other hand, DVC values calculated using nonlinear decoders such as kernel SVM and multi-layer perceptron are significantly lower, suggesting overfitting (Fig. A.3e).
Figure A.3: The main results are robust to particular choices in the implementation. We tested the result in Fig. 2c where networks with better ImageNet-1k performance are less alignment under different conditions when (a) using the top 10 principal components (PCs) instead of 25. (b) Using the top 50 PCs instead of 25. (c) Using Spearman’s rank correlation instead of Pearson’s correlation. (d) Using an eigen SVD solver with shrinkage instead of a SVD solver for LDA. (e) Measure of monkey-monkey DVC is consistent across linear decoders and across random splits. Error bar indicates the standard deviation.
C.4 DVC results are consistent in another neural datasets
To alleviate concern that the dataset from which we derived the main results contains only two monkeys, we verified the result on an additional dataset from Bashivan et al [Bashivan2019neural]. The dataset contains three monkey subjects viewing natural and synthesized images. Among the three subjects, only one (’monkey M’) has sufficient neuron count for this test (n = 168). The neurons are pooled across four recording sessions. We used data from monkey M viewing naturalistic images and calculated its DVC against neural networks viewing the same images. The main results are also significant in this case (Pearson’s correlation = -0.84, p = 0.87e-05, Fig.A.4a).
Figure A.4: The main result is confirmed in an additional dataset from [Bashivan2019neural]. (a) DVC of monkey M against representative models. Again, these is a strong negative correlation between the top-1 accuracy of the deep networks and the DVC to the monkey brain. (b) DVC between representative models.
C.5 DVC between representations and raw image pixels
We calculated the DVC between the V4/IT representations and model hidden representations to flattened image pixels and found that the DVC between the two monkeys and raw image pixels are 0.153 and 0.195, respectively. The DVC between model hidden representations and network DVC are generally lower but still substantial (Fig. A.5a)). The results suggest that part of the DVC results might come from remnants of linear separability in the image pixels, which is reduced by nonlinearity in the brain and models. Indeed, the LDA score for categorization using raw image pixels is 0.70, which is smaller than that of brain and model representations (>0.90) but still very high. We plotted the LDA axis for classifying ’chair’ and ’car’ in image pixel space and see clear silouettes of the objects (Fig. A.5b), suggesting that high DVC to image pixels may be due to simplicity of generated image samples. It would be interesting to test the result on datasets with more complex and diverse images.
Figure A.5: DVC between the monkey and model representations and raw image pixels. (a) The DVC between models and raw image pixels are markedly lower than DVC between models and monkeys yet still substantial. (b) The LDA axis between categories ’chair’ and ’car’ in image pixel space.
C.6 Split normalization under uncommon conditions
@@
The split normalization procedure could behave counterintuitively under extreme conditions. For example, it could result in a DVC value larger than 1 when the internal noise is high. None of the normalized correlations between different representations reached the cap. In addition, we took the absolute value before calculating the geometric means. When there is no correlation between two representations, this would bias the normalized DVC to a small positive value. None of the representations in this study fall in this range.
@@
Appendix D Additional results on the effect of behavioral decoding on Cohen’s Kappa
@@
Geirhos et al. [geirhosAccuracyQuantifyingTrialbytrial2020] have discussed several caveats of applying Cohen’s Kappa, including that Cohen’s Kappa is bounded by error overlap expected by chance $c_{exp}$. We also explicitly showed that the accuracy difference $d=|p_{i}-p_{j}|$ provides an upper bound on Cohen’s Kappa (see Section A.2). When the accuracy difference is high, one subject is often right, while the other is often wrong, then their behaviors cannot be consistent. From the original behavioral data, it is clear that human subjects perform with high accuracy but models perform poorly(Fig. A.6). While Geirhos et al used simulations to show that under the condition that the subjects act independently, accuracy is not necessarily correlated with model performance, Cohen’s Kappa may still depend on accuracy under more general conditions.
Figure A.6: Re-plotting of the data (based on "edge" stimuli from [geirhosAccuracyQuantifyingTrialbytrial2020]. Red: accuracy of human subjects. Blue: accuracy of deep network models tested. Overall, the accuracy of the deep network models are much lower than that of human subjects.
@@
We would like to verify whether the high model-model consistency reported in the original work might be inflated by the low-accuracy high-bias condition caused by the choice of decoder (directly aggregating the probabilities). To test this, we take the original stimuli provided by Geirhos et al., and calculate Cohen’s Kappa between the models by (i) taking the average of the probability of the corresponding ImageNet-1k subclasses or (ii) training a 5-fold cross-validated logistic regression decoder on the representations in the penultimate layer. The result shows that compared to (i), approach (ii) achieves higher accuracy (Fig. A.7b,c), exhibits less bias towards certain categories (Fig. A.7d,e) and results in significantly lower model-model consistency as measured by Cohen’s Kappa (Fig. A.7a).
@@
Figure A.7: Using logistic regression decoder results in higher accuracy, lower bias and lower Cohen’s Kappa estimate compared to mean probability decoder, estimated on the ’edges’ images. (a) Cohen’s Kappa estimated using a mean probability decoder is significantly higher than that estimated by a logistic regression decoder. (b) Behavioral accuracy of mean probability decoder. (c) Behavioral accuracy of logistic regression decoder. (d) Choice histogram of the mean probabilities decoder. (e) Choice histogram of the logistic regression decoder.
@@
While the shared behavioral bias that results from aggregating probabilities from the original ImageNet-1k classes is very interesting, it does make the human-model consistency and the model-model consistency a lot more ambiguous. Thus, it may be more appropriate to use a stronger decoder for classification or to use DVC when applicable.
Appendix E Comparison to Representational Similarity Analysis
Representational similarity analysis (RSA) measures the similarity between the general arrangement of categories (or samples) between two representations. Under appropriate conditions, it can be seen as equivalent to centered kernel alignment (CKA) and canonical correlation analysis (CCA) [williamsEquivalenceRepresentationalSimilarity2024]. In the monkey dataset that we tested, RSA is positively correlated with DVC when measuring both model-monkey alignments and model-model alignments (see Fig. A.8a-d)).
While RSA may contain information about the correlations of task-related dimensions, they may be confounded by task-irrelevant correlations. We demonstrate this point using a simple simulation. Given the same baseline decision variables (’task-relevant correlation’), we introduced independent fluctuations (’noise’) and shared fluctuations (’task-irrelevant correlation’). We find that while DVC is generally unaffected by the introduction of stronger task-irrelevant correlation (Fig. A.8e). In contrast, RSA quickly plateaus as task-irrelevant correlation overshadows task-relevant ones. This simulation demonstrates that RSA and DVC capture different aspects of the representation. The ability of DVC to focus on only task-relevant dimensions makes it suitable when applied to scenarios with high shared fluctuation based on common stimulus input, where only some dimensions encode task-relevant information.
Figure A.8: Comparison to Representational Similarity Analysis (RSA). (a) Heatmaps showing the DVC and representational similarity (RS) for pairs of representative models. (b) There is a strong positive correlation between DVC and RS on model-model consistency (evaluated on this dataset). (c) The correlation between category-level RSA. (d) Category-level RSA recreates the main result that alignment between model and brain declines with ImageNet top-1 accuracy. (e) Simulation results show that RSA is conflated with task-irrelevant correlations whereas DVC only cares about task-relevant ones.
Appendix F Broader Impacts
@@
We expect DVC to be a broadly applicable approach to study the similarity of brain and neural network models. On the positive side, the method and results discussed in this paper could redirect community focus away from brute-force scaling and toward more targeted investigations into task-relevant representation alignment and model-brain convergence. It could in the long term lead to models that are more brain-like, thus greatly facilitating research in fields like neuroscience, cognitive science and AI interpretability and safety. However, while DVC offers a biologically grounded lens for comparing model and brain representations, promoting alignment with biological brains might inadvertently constrain models in certain domains where brain cognition is suboptimal.